E and D Time - 20140614 UT

Observers: O. Kuhn, J. Hill, D. Miller (Remote from Tucson)
Telescope Operator: S. Allanson (LBTO)
AO Support: D. Miller (LBTO)
Telescope Support: M. Midkiff (LBTO)
Software Support: T. Golota, C. Biddick (Remote from Tucson) D. Summers (LBTO)
Instrument Support: J. Power (LBTO)

Summary

Finished the testing of and released Build 2014C including use with adaptive optics. Doug S did a lot of testing on the alarm handler.

Got more non-sidereal testing done. For LUCI1 (and by implications LUCI2 and MODS2), it appears to be working OK, and is now ready for use on a real asteroid when LUCI is working. For MODS1, the probe offset to the guiding hotspot is still giving us fits. Hopefully we did enough sample presets for Taras to track down the bug.

Issue 5212 (intermittent slow guide update response from PCS) could limit the rate at which non-sidereal guiding can be used. It was already giving some trouble at elevations above 85 deg where guide corrections can be large. This is not a new problem in 2014C.

Olga and John made more progress with testing WRS on-sky using LBC-Red in parallel with other activities. Next thing to be done is to rotate the Zernikes from the observed frame to the frame of the primary mirror. JMH did this the following evening.. WRS is still having problems getting the amplitude of the coma correct when the coma is small.

Doug M did more work on the Z4/Z11 crosstalk. He suspects a bug in PSF (see 7:20 UT below)

Doug M did various tests with AO Tip-Tilt mode and LUCI1 including setting the focus of "W" for LUCI1. LUCI1 was OK for this activity, but much too noisy for deep science.

Details

TCS 2014C Build

Chris and Taras ran tests of Build 2014C in late afternoon and early evening.

Open

MODS1 in pseudomono mode

Running GCS v2.17.0

02:36 Open clear sky with the wind just below the limit.

02:49 ACQUIRE preset - default collimation is outstanding

BS9158

02:53 ACTIVE preset sidereal - acquisition fails from background artifacts

02:55 ACTIVE preset sidereal - succeeds (still a bit bright for wfs)

03:10 ACTIVE preset NS with zero rate

03:14 ACTIVE preset NS -- succeeds briefly, but then probe moves back to 0,0

03:21 ACTIVE preset NS -- succeeds briefly, but then probe moves back to 0,0

03:25 ACTIVE preset NS w/ DEC rate -- (03:25:27 hangs with IIF talking to PCS)

03:29 Kill PCS then had to do end of night to unstick the preset (Chris got to see it.)

03:39 ACTIVE preset NS w/ DEC rate -- succeeds briefly, but then probe moves back to 0,0

03:42 ACTIVE preset sidereal gs=5 -- succeeds but has a problem with 8 second intervals in the guide stabilization images.

WRS failed - near 04:00 with a saturated star in the image, but not the whole image saturated. lbcr.20140614.040236.

04:00 Stop and restart GCSL and GCSR to run v2.17.1

selectAGW left MODS_L.cfg

04:06 ACTIVE preset NS -- succeeds briefly, but then probe moves back to 0,0 (are we really running the new version?)

04:14 Stop and restart GCSL and GCSR to run v2.17.2

selectAGW left MODS_L.cfg

still has timeout on MODS power status

04:17 ACTIVE preset NS -- succeeds briefly, but then probe moves back to 0,0

04:18 Stop and restart GCSL and GCSR to run v2.17.0

selectAGW left MODS_L.cfg

LUCI1 NS Tests

04:22 switch to LUCI1 on left

04:29 Kill and restart AOSR

Authorize failure.....

04:33 Stop AOSR and leave it off

04:33 Authorize LUCI and LBC again - successful

04:37 ACQUIRE preset has long delay

04:40 ACTIVE preset sidereal

04:42 guide jump on SX - coma correction ??

04:45 offset -- guide star position missed by ~1 arcsec -- because we at EL=85.5 ??

04:46 offset -- guide star position missed by ~0.6 arcsec

04:48 ACTIVE preset NS w/zero rate -- guiding is still slowed by moving the probe after each guide iteration

04:53 ACTIVE preset NS --

2014-06-14T04:52:49.631400+00:00 tcs4 LBT_left_GCS: GT: eff.GuideRate: 0.033 Hz max.GuideRate: 0.033 Hz expTime: 2.000 s time4PCSxfer: 18.107 s time4ImgAcq: 2.653 s processing overhead: 9.325 s
This became Issue 5212.

Problems from probe motion uncertainty ???

05:xx switch to off-axis guide star so we can take LUCI images on-axis

(details from log missing in this area)

05:30 timeout for pause guiding for an offset is too short given all these other delays

05:34 Series of 30 x 4 sec LUCI images starts with 0018 (RA rate is +30 dRA)

05:43 ACTIVE preset NS dDec +30

05:45 Series of 30 x 4 sec LUCI images starts with 0048 (RA rate is +30 dDec)

Telescope moved in the proper direction

Crosstalk Coefficients

05:52 ACTIVE preset off-axis sidereal (probe couldn't reach guide star)

05:55 ACTIVE preset off-axis sidereal

06:15 Set all crosstalk factors to 0.0 in psf.conf in order to measure new Z11->z4, Z11->Z22, Z22->Z4 and Z22->Z11
  • /home/dlmiller/data/20140614/zern_gain_4_11_22
  • Z4 in sext_4_22_zern.log
    • No crosstalk to other modes
  • Z11 and Z22 in z4_11_22_zern.log
    • Z11->Z4 = 4.5
    • Z11->Z22 = 0.0
    • Z22->Z4 = 3.5
    • Z22->Z11 = -1.1

06:51 Set all crosstalk values to the values I calculated and tested before, with the exception fo z11->z4 set to 4.25 (old value)
  • /home/dlmiller/data/20140614/zern_gain_12_13_more

07:20 Set ih psf.conf
    • Z11->Z4 = 4.5
    • Z11->Z22 = 0.0
    • Z22->Z4 = 3.5
    • Z22->Z11 = -1.1
    • However, now measure Z11 of 100 rather then 500 that was sent.
    • Similar for Z22
    • Error in Chris' routine?

AOTT with LUCI1

07:30 iif_preset, GC_M13, MODE=ACE, GS=6, TS=76, RS=76

07:45 First image luci 81, DIT=4.0, NDIT=1

07:50 Offset 5,0 = 82

07:53 iif_preset, GC_M13, MODE=ACE, RS=6, TS=6, GS=76, PA=180

07:57 Luci 83

07:58 iif_runao

07:59 Low order (10 modes) closed loop, LUCI 84

08:01 Closed Loop, bin 2, 120 modes, 625 Hz. Luci 85
  • mag estimate = 12.1

08:03 DIT=4.0, NDIT=5, Luci 86

08:06 iif_offset, 120, 0 (guide probe error)

08:09 Luci 87: FWHM 6.2, 5.7 z = 44.5

08:17 Luci 88: FWHM 5.5, 5.3 Z=45.5

08:20 Luci 89: FWHM 5.6, 4.9 Z=46.5

08:24 Luci 90: FWHM 5.3, 4.5 Z=47.5

08:28 Luci 91: FWHM 5.5, 4.6 Z=48.5

08:30 Luci 92: FWHM 5.5, 4.5 Z=49.5

08:32 Luci 93: FWHM 4.8, 3.7 Z=50.5

08:35 Luci 94: FWHM 4.9, 3.6 Z=51.5

08:36 Luci 95: FWHM 4.2, 3.3 Z=52.5

08:39 Luci 96: FWHM 3.6, 2.7 Z=53.5

08:42 Luci 97: FWHM 3.5, 2.6 Z=54.5

08:45 Luci 98: FWHM 3.3, 2.6 Z=55.5

08:49 Luci 99: FWHM 3.4, 2.8 Z=56.0

08:50 Turned off software limit on Z-Stage

08:49 Luci 100: FWHM 3.8, 2.9 Z=56.5

08:52 Luci 101: FWHM 3.6, 3.3 Z=57.0

08:55 Luci 102: FWHM 3.7, 3.6 Z=58.0

08:56 Luci 103: FWHM 4.4, 4.1 Z=59.0

08:49 Luci 104: FWHM 3.4, 2.6 Z=54.5

09:00 Pause AO on engineering GUI

09:02 Luci 105: FWHM 5.8, 3.8 Z=54.5 (at this value from here on out)

09:04 Luci 106: FWHM 6.2, 4.1

09:09 Center guide star on guider hotspot
  • x-stage -77.65 (less negative moves star to the left)
  • y-stage -51.4 (more negative moves star down)
  • Z4 on WFSC = -100 nm
  • Z5 = 140 nm
  • z6 = -200
  • z7 = -100
  • Z8 = -80
  • z9 = -15
  • z10 = 24
  • z11 = -25
  • wfscimage 303

09:14 iif_offset, 30 , 0 to center the WFSC pupil

09:16 AO Open Loop, Luci 107: FWHM 4.3, 3.9 Almost as good as AO Closed loop!!

09:20 AO Open Loop, Luci 108: FWHM 4.4, 4.0 Almost as good as AO Closed loop!!

09:23 iif_preset, GC_M13, MODE=ACE, RS=6, TS=6, GS=76, PA=180: Failed, Steve change pointing

09:26 iif_preset, GC_M13, MODE=ACE, RS=6, TS=6, GS=76, PA=180: Succeed

09:29 iif_runao
  • mag estimate= 12.2

09:40 AO Loop Closed: Luci 109: FWHM 5.8, 6.3, Z = 44.5

09:48 AO Loop Closed: Luci 110: FWHM 3.4, 2.6, Z = 54.5

09:51 irc_pauseao

09:51 AO Loop Open: Luci 111: FWHM 4.8, 4.4

09:52 irc_resumeao

09:52 AO Loop Closed: Luci 110: FWHM 4.0, 3.0, Z = 54.5

10:32 iif_preset, GC_M39, MODE=ACE, RS=75, TS=75, GS=59, PA=180: Succeed, iif_runao
  • Still bin 2

10:34 iif_preset, AO745, MODE=ACE, RS=0, TS=0, GS=23, PA=180. Bin 1, 900 hz, 400 modes

10:56 AO Loop Closed: Luci 113: FWHM 6.0, 5.5, Z = 44.5

11:02 AO Loop Closed: Luci 114: FWHM 2.2, 2.1, Z = 54.5

11:05 AO Loop Open: Luci 115: FWHM 5.3, 4.3

11:07 AO Loop Closed: Luci 116: FWHM 1.6, 1.6, Z = 54.5

11:09 AO Loop Open: Luci 117: FWHM 4.5, 4.5

11:20 iif_preset, OC_NGC1027, MODE=ACE, RS=32, TS=32, GS=78, PA=180. Bin 2, 400 hz, 400 modes

11:25 AO Loop Open: Luci 118: FWHM 9.6, 9.1

11:33 Too bright, could not close AO loop. Safe Skip

11:35 Steve Close Up

-- %USERSIG{JohnHill - 2014-06-14}%
Topic revision: r5 - 15 Jun 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback